# Smart Exploration in Reinforcement Learning Using Bounded Uncertainty Models

### v2

@@ VI CONCLUSIONS AND FUTURE WORK
In this paper, we proposed an exploration strategy for reinforcement learning that leverages model-based Q-function bounds to guide the agent’s exploration. Our work establishes multiple theoretical results that guarantee the convergence of the proposed exploration strategy to the optimal policy in a general MDP setting. Furthermore, in the finite state and action space case, and under reasonable assumptions on the model set, we obtain a practical algorithm, BUMEX, with the same convergence guarantees. We also demonstrated that, under additional assumptions on the transition probabilities, our method achieves finite-time convergence to the optimal policy. The effectiveness of the proposed strategy was validated in several Gym environments, where it outperformed standard exploration methods.
Future work includes extending the theoretical results to infinite state-action spaces, e.g., through function approximation. Additionally, a sample complexity analysis of BUMEX is desired to quantify the efficiency compared to state-of-the-art RL methods.

### v3

@@ VI CONCLUSIONS AND FUTURE WORKS
In this paper, we proposed a novel exploration strategy for reinforcement learning that leverages model-based Q-function bounds to guide the agent’s exploration. Our work establishes multiple theoretical results that guarantee the convergence of the proposed exploration strategy to the optimal policy in a general MDP setting. Furthermore, in the finite state and action space case, and under reasonable assumptions on the model set, we obtain a practical algorithm with the same convergence guarantees. We also demonstrated that, under additional assumptions on the transition probabilities, our method achieves finite-time convergence to the optimal policy. The effectiveness of the proposed strategy was validated in several gym environments, where it outperformed standard exploration methods.
For future work, we plan to apply the practical algorithm to a broader range of examples and to extend the theoretical results to derive a practical algorithm for the infinite-state, infinite-action case.